OpenStudy (anonymous):

Is this mapping linear? L(x1, x2) = (0, x1 + 2, x2)

OpenStudy (anonymous):

no its not linear

OpenStudy (anonymous):

Thanks, how can you tell?

OpenStudy (perl):

Use the definition of linear transformation L(x + y) = L(x) + L(y) L ( c*x) = c*L(x) , where x,y are vectors

OpenStudy (perl):

c is a scalar

OpenStudy (perl):

If you suspect this function is not a linear transformation, you just need to find a specific example that demonstrates the failure of one of the properties.

OpenStudy (perl):

how about try x = (1,2) y= (3,4) Now show that L( x + y ) =/= L(x) + L(y) , and you have produced a counterexample.

OpenStudy (perl):

in other words show that L( (1+3, 2 + 4) ) =/= L((1,2)) + L((3,4))

OpenStudy (perl):

another way to prove it, show that L( 2*(1,1)) =/= 2 * L(1,1) this might be shorter

OpenStudy (anonymous):

Thanks, got it